### v1

# Angle-dependent planar thermal Hall effect by quasi-ballistic phonons in black phosphorus

Abstract
The origin of the phonon thermal Hall effect in insulators is a matter of ongoing debate. The large amplitude of the signal in an elemental non-magnetic solid, such as black phosphorus (BP) calls for a minimal mechanism with no role for spin degree of freedom. Here, we show that a longitudinal heat flow generates a transverse temperature gradient in BP even when the magnetic field, the heat current and the thermal gradient lie in the same plane. The long phonon mean-free-path leaves little room for scattering by point-like symmetry breaking defects. we show that the signal peaks when the magnetic field is oriented along one of the two diagonal orientations of the puckered honeycomb plane and argue that this can be understood as the sum of two distinct contributions each parallel to a mirror plane. This angular dependence as well as the order of magnitude of the observed signal points to the torque exerted by magnetic field on electric dipoles traveling with heat-carrying phonons as the driver of the effect.
The thermal conductivity is a second rank tensor linking the heat current density and the temperature gradient vectors. The thermal Hall effect (THE) refers to non-zero off-diagonal components of this tensor, $\overline{\kappa}$. Its phonon-driven origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26] interest. More recently, attention has focused on the observation of a planar thermal Hall effect [16, 17, 15]. This is a configuration with the three relevant vectors (Heat current, temperature gradient and magnetic field) lying in the same plane. The persistence of the signal when the signal is forbidden by crystal symmetry was attributed to defects breaking the local symmetry [17].
Figure 1: Crystal structure, phonon mean free path and two versions of the thermal Hall effect.(a) Top (two layers) and side (single layer) view of the lattice of black phosphorus (BP). P atoms, marked in blue and red, belong to atomic sites with distinct environments. The $x$ and $z$ axes respectively correspond to the armchair and zigzag directions of the BP layer plane. (b) Longitudinal thermal conductivity (from previous work [12]) along the armchair ($\kappa_{xx}$) and the zigzag ($\kappa_{zz}$) orientation, at least four times of anisotropy can be seen. (c) The temperature dependence of mean free path of phonon ($l_{ph}$) in BP for two in-plane orientations ($J_{q}||z$ axis and $J_{q}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat. It shows an increasing behavior before 10 K with cooling and is comparable to the sample thickness around 20 K, the peak temperature of $\kappa_{ii}$. (d) The schematic for measuring the thermal Hall effect, with the magnetic field ($B$) perpendicular to the longitudinal heat current ($J_{q}$) and the transverse thermal gradient ($(\nabla T)_{\perp}$). The thermal Hall conductivity tensor is denoted as $\kappa_{ij}^{k}$. (e) The schematic for measuring planar thermal Hall effect, with the magnetic field ($B$) rotating from the orthogonal to the parallel direction of the $J_{q}$ or the $(\nabla T)_{\perp}$. Two types of planar thermal Hall conductivity tensors are denoted as $\kappa_{ij}^{i}$ and $\kappa_{ij}^{j}$ respectively.
Figure 2: Planar thermal Hall data. (a) The field dependence of thermal Hall angle $\nabla T_{j}/\nabla T_{i}$ , with $J_{q}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures. This allows to quantify $\kappa_{xz}^{x}$. (b) Same for $J_{q}$ along the $x$ axis, $B$ and $(\nabla T)_{\perp}$ along the $z$ axis. This allows to quantify $\kappa_{xz}^{z}$. (c) Same for $J_{q}$ and $B$ along the $z$ axis and $(\nabla T)_{\perp}$ along the $x$ axis. This allows to quantify $\kappa_{zx}^{z}$. (d) The temperature dependence of the thermal Hall angle for three different configurations at 12 T. All measurements are performed on the same sample. Note that the thermal Hall angle, $\nabla T_{j}/\nabla T_{i}$ is almost unchanged when the orientations of $B$ rotates from $J_{q}$ to $(\nabla T)_{\perp}$. On the other hand, it changes by a factor of six times when the orientations of $J_{q}$ and $(\nabla T)_{\perp}$ permute with each other. This difference combined with the anisotropy in longitudinal thermal conductivity warrants Onsager reciprocity: $\kappa_{xz}^{z}=-\kappa_{zx}^{z}$, as seen in (e). (f) Comparison of thermal Hall conductivity $\kappa_{xz}$ with the magnetic field $B$ rotating from the orthogonal ($y$ axis, from previous work [12] with the same sample) to the parallel direction of the $J_{q}$ ($x$ axis) or the $(\nabla T)_{\perp}$ ($z$ axis).
[…]
Figure 4: From atomic displacements to dipolar moments (a) Lattice distortion due to the thermal expansion. (b) Phosphorus atoms surrounding an inversion center. Their displacements associated with an acoustic mode (top) will affect the local electric dipole moments (middle), generating a net dipole moment (bottom). (c) Phonon modes are associated with waves of dipolar moments. A static magnetic field will exert a torque on the dipole moment traveling with the sound wave. (d) This torque ($\tau=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$) will survive only when the magnetic field is oriented perpendicular to the wave velocity. Orange arrows show the orientation of the Lorenz forces on the two poles of the dipole.
Here, we report on the observation of a planar Hall effect in black phosphorus(P), an elemental insulator with ballistic phonons [27] and no magnetism. The thermal conductivity tensor has unequal diagonal components ($\kappa_{zz}\simeq 5\kappa_{xx}$). Nevertheless, the off-diagonal components match each other ($\kappa_{xz}(B)\simeq\kappa_{zx}(-B)$), as expected by Onsager reciprocity. We quantify the variation of the thermal Hall signal as a function of the in-plane orientation of the magnetic field and find that it displays a twofold oscillation with minimum and maximum along one of the two diagonals of the $xz$ plane of the puckered honeycomb lattice. This indicates that the signal is the sum of two sinusoidal contributions along two high-symmetry axes. A quantitative account of our observation is missing. Nevertheless, we argue that a scenario in which electric dipole moments, generated by the thermal gradient and the atomic vibrations, play a key role is compatible with the order of magnitude of the observed signal as well as its angular variation.
Fig. 1a shows the crystal structure of black P. Identical phosphorus atoms located on two distinct sites are marked in blue and red. Each layer is a puckered honeycomb lattice in the $xz$ crystallographic plane, where the $x$ and $z$ axes correspond to the armchair and zigzag orientations. As seen in Fig. 1b, the longitudinal thermal conductivity, as found previously [27, 28], is significantly different along the two orientations. Along the zigzag orientation, thermal conductivity is much larger than along the armchair orientation in the low temperature limit. The sound velocity shows a similar but attenuated anisotropy (9.6 km/s along $z$-axis vs. 4.6 km/s along the $x$-axis) [29]. As a consequence, the phonon mean free path $\ell_{ph}(T)$ , shown in Fig. 1c, is also anisotropic. It is twice longer along the z-axis and approaches the sample thickness (30 $\mu$m) around 20 K. Thus, at this temperature phonons become quasi-ballistic and accordingly thermal conductivity becomes size dependent [27].
A Hall response refers to a signal odd in magnetic field and oriented perpendicular to the applied current. Designating the transverse temperature gradient by $(\nabla T)_{\perp}$, the applied longitudinal heat current by $J_{q}$ and the magnetic field by $B$, the standard thermal Hall configuration corresponds to $B\perp J_{q}\perp(\nabla T)_{\perp}$. It is sketched in panel d and was the subject of our previous study [12]. Panel e shows the configuration for planar thermal Hall effect, which corresponds to $B~{}||~{}(J_{q}\perp(\nabla T)_{\perp})$. We use $\kappa_{ij}^{k}$ to designate the $\overline{\kappa}$ component corresponding to $J_{q}$ along the $i$-axis, $(\nabla T)_{\perp}$ along the $j$-axis and the magnetic field along the $k$-axis.
Fig. 2a-c shows the planar thermal Hall data for three different configurations. In each panel, the orientations of the three relevant vectors (B, $J_{q}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla T_{j}/\nabla T_{i}$) is shown.
Fig. 2d compares the temperature dependence of the thermal Hall angle in three planar configurations at 12 T. It is finite in the three planar configurations, but there is a large difference in amplitude when one permutes the orientations of $J_{q}$ and $(\nabla T)_{\perp}$. Since the longitudinal thermal conductivity is anisotropic by the same factor, this difference warrants an equality between the absolute values of $\kappa_{xz}^{z}$ and $\kappa_{zx}^{z}$, as expected by Onsager reciprocity.
The thermal Hall angle (Fig. 2d) combined with the longitudinal thermal conductivity (Fig. 1b) leads to the planar thermal Hall conductivity, which is shown in Fig. 2e and f. The thermal Hall response is finite in four distinct configurations. We can see not only the validity of the Onsager reciprocity (Fig. 2e), but also the very small influence of the orientation of magnetic field on the amplitude of $\kappa_{xz}$ (Fig. 2f).
[…]
Thus, we detect in a crystal belonging with the $mmm$ point group symmetry, there is a planar thermal Hall signal with bi-axial symmetry. Let us note that at peak temperature of this thermal Hall signal, the longitudinal thermal conductivity of black P is as large as $\kappa_{p}\simeq 2000$ W/Km. This is to be compared with cases such as $\alpha$-RuCl3 ($\kappa_{p}\simeq$ 3 W/Km) [11], Na2Co2TeO6 ($\kappa_{p}\simeq$ 10 W/Km) [16] or Y-kapellasite ($\kappa_{p}\simeq$ 2 W/Km) [17]. Contrary to these cases, the phonon mean free path in black P is close to the sample dimensions and local defects are unlikely to play a major role.
In anisotropic dense medium, thermal diffusivity is a tensor, $\overline{D}$. The heat equation becomes :
$\frac{\partial T}{\partial t}=\overrightarrow{\nabla}\cdot(\overline{D}\overrightarrow{\nabla}T)$ (1)
Consider now the energy continuity equation, which relates the energy flux to specific heat per volume, $C$:
$\overrightarrow{\nabla}\cdot\overrightarrow{J_{q}}+C\frac{\partial T}{\partial t}=0$ (2)
The combination of the two previous equations leads to :
$\overrightarrow{J_{q}}=-C\overline{D}\overrightarrow{\nabla}T$ (3)
The thermal conductivity tensor $\overline{\kappa}$, is thus the thermal diffusivity tensor multiplied by a scalar: $\overline{\kappa}=\overline{D}C$. Therefore, off-diagonal components in $\overline{\kappa}$ are proportional to their counterparts in $\overline{D}$. The latter is the product of carrier velocity and carrier mean free path. Sound velocity, $v_{s}$, and phonon mean free path $\ell_{ph}$ (see Fig. 1c) are both anisotropic at zero magnetic field. The issue is to find a way for magnetic field to skew one or both of these tensors off the symmetry axes
Recent theoretical studies have scrutinized intrinsic phonon band[31] incorporated with energy magnetization contribution [32, 21] and phonon angular momentum acquired through interaction with the spins [33, 34] or with orbital motion of ions [35].
Absent magnetism, the interplay between electric polarization and phonons deserves scrutiny. The computed map of charge distribution is strongly orientational [36]. Dipole-active phonon modes has been detected by infrared spectroscopy [37]. These are optical phonons and not heat-carrying acoustic modes which interest us. Nevertheless, any phonon breaking the inversion symmetry will generate an electric dipolar wave. This is a feature highlight in the context of ferrons, the elementary excitation of a ferroelectric solid, even in a paraelectric insulators [38, 39].
The thermal expansion of black P is not only anisotropic, but has opposite signs along armchair and zigzag crystalline orientations [40]. As a consequence, in presence of a temperature gradient each tetragonal unit cell is slightly distorted. Its lattice parameter, compared to its colder or warmer neighboring cells, is longer and shorter along opposite (Fig. 4a) orientations. This paves the way for finite dipolar polarization, and for chirality in presence of the magnetic field. The inversion center of the pristine unit cell is not an atomic site and atomic displacements associated with phonons generate electric dipole moments (Fig. 4b). Therefore, acoustic phonons breaking inversion centers are traveling electric dipolar moments capable of coupling to a static magnetic field (Fig. 4c).
The torque exerted by a static magnetic field, $\overrightarrow{B}$, on an electric dipole of $\overrightarrow{P}$ moving with a velocity of $\overrightarrow{v}$ is : $\overrightarrow{\tau}=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$ [41]. Let us rewrite it as:
$\overrightarrow{\tau}=(\overrightarrow{P}\cdot\overrightarrow{v})\overrightarrow{B}-(\overrightarrow{P}\cdot\overrightarrow{B})\overrightarrow{v}$ (4)
Thus, the torque due to the Lorenz force exerted on each of the two poles is finite when the magnetic field is oriented perpendicular to the orientation of the dipole propagation (Fig. 4d). These are ingredients for a scenario for planar thermal Hall signal in a non-magnetic insulator (see the supplement [30] for a discussion of the expected angular variation). A rigorous treatment is beyond the scope of this paper. However, let us compare the order of magnitude of the expected and the measured signals.
Experimentally, the thermal Hall angle at $B$ = 10 T peaks to ×$\approx\times 10^{-2}$ in black P. The length scale extracted from this angle and fundamental constants ($\lambda_{tha}$ = $\ell_{B}\sqrt{\kappa_{ij}/\kappa_{jj}}$) is about 5 Å. Intriguingly, in all insulators displaying a thermal Hall signal, this length scale lies in a narrow range (2 Å${}<\lambda_{tha}<7$ [12, 14]. As seen in figure S2 in the supplement [30], this remains true for recently studied cases of Si, Ge and quartz [15].
Suppose an electric dipole moment of $\delta p$ traveling with the velocity of sound, $v_{s}$ (loosely linked to the Debye frequency by $v_{s}\sim a\omega_{D}$, where $a$ is the interatomic distance). The exerted torque, that is the angular derivative of the energy, would be $Ba\omega_{D}\delta p$. Divided by the Debye energy, it yields $\theta_{H}\simeq\frac{a\delta p}{\hbar}B$, the order of magnitude of the expected rotation angle. With $\ell_{B}^{2}=\frac{\hbar}{eB}$, assuming $\delta p\approx ea$, leads to:
$\theta_{H}\approx\frac{a^{2}}{\ell_{B}^{2}}$ (5)
This expectation is of the order of magnitude of the measured signal. A rigorous treatment would presumably include in this expression the Grüneisen parameter [42]. Possibly, the atomic electric dipole polarizability [43] would also be present. However, the former, which is dimensionless remains of the order of unity and the latter introduces a length scale which does not vary widely among different solids. Therefore, a scenario attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves emerges promising from our study.
We thank Benoît Fauqué for stimulating discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2022YFA1403500), the National Science Foundation of China (Grant No. 12004123, 51861135104 and 11574097 ), and the Fundamental Research Funds for the Central Universities (Grant no. 2019kfyXMBZ071). X. L. was supported by The National Key Research and Development Program of China (Grant No.2023YFA1609600) and the National Science Foundation of China (Grant No. 12304065).
Supplementary Materials for “Angle-dependent planar thermal Hall effect in black phosphorus”
S1 Materials and Methods
Black phosphorus crystals used in this work are same to our previous report [34]. They were synthesized under high pressure and came from two different sources. Sample #1 was obtained commercially and sample #2 was kindly provided by Prof. Yuichi Akahama (University of Hyogo).
[…]
The longitudinal ($\nabla T_{i}$) and the transverse ($\nabla T_{j}$ ) thermal gradient generated by a longitudinal thermal current $J_{q}$ were measured. They lead to the longitudinal ($\kappa_{ii}$) and the transverse ($\kappa_{ij}$) thermal conductivity, as well as the thermal Hall angle ($\nabla T_{j}/\nabla T_{i}$):
$\kappa_{ii}=\frac{Q_{i}}{\nabla T_{i}}$ (S1)
$\frac{\nabla T_{j}}{\nabla T_{i}}=\frac{\kappa_{ij}}{\kappa_{jj}}$ (S2)
$\kappa_{ij}=\frac{\nabla T_{j}}{\nabla T_{i}}\cdot\kappa_{jj}$ (S3)
Here $Q$ is the heat power.
S2 Raw data of angular dependent planar thermal Hall effect
Fig. S1 shows the field dependent thermal Hall angles ($\nabla T_{j}/\nabla T_{i}$) with the $J_{q}$ alongs $x$ axis and the $(\nabla T)_{\perp}$ along $z$ axis at six different field orientations. $\phi$ is the angle of the magnetic field with respect to the $J_{q}$ ($x$ axis). $\phi=0^{\circ}$ and $\phi=90^{\circ}$ represent the orientations along two high-symmetry axes, and their $\nabla T_{j}/\nabla T_{i}$ signal have the same amplitude but opposite signs. $\phi=-45^{\circ}$ and $\phi=45^{\circ}$ represent two diagonal orientations of the puckered honeycomb plane that have the maximum and vanished responses respectively, suggesting that the planar thermal Hall signal consists of two distinct contributions each along one high-symmetry axis. $\phi=-45^{\circ}$ and $\phi=135^{\circ}$ have signals with opposite signs, as expected from the antisymmetric operation ($\kappa_{ij}$(B) = -$\kappa_{ij}$(-B)).
Figure S1: Raw data of angular dependent planar thermal Hall effect in Black phosphorus. Field dependent thermal Hall angle ($\nabla T_{z}/\nabla T_{x}$) at a constant temperature and six different field orientations.
Figure S2: Thermal Hall angle in different insulators The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source:  [1, 2, 3, 4, 5, 6, 8, 9, 11, 12, 14, 15]).
S3 Torque exerted by the magnetic field on moving dipoles
The torque exerted by a static magnetic field, $\overrightarrow{B}$, on an electric dipole of $\overrightarrow{P}$ moving with a velocity of $\overrightarrow{v}$ is :
$\begin{split}\overrightarrow{\tau}&=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})\\ &=(\overrightarrow{v}\times\overrightarrow{B})\times\overrightarrow{P}\\ &=(\overrightarrow{P}\cdot\overrightarrow{v})\overrightarrow{B}-(\overrightarrow{P}\cdot\overrightarrow{B})\overrightarrow{v}\end{split}$ (S4)
The electric dipole $\overrightarrow{P}$, field $\overrightarrow{B}$, and the velocity $\overrightarrow{v}$ can be expanded as:
$\begin{split}\overrightarrow{P}=P_{x}\overrightarrow{e_{x}}+P_{y}\overrightarrow{e_{y}}+P_{z}\overrightarrow{e_{z}};\\ \\ \overrightarrow{B}=B_{x}\overrightarrow{e_{x}}+B_{y}\overrightarrow{e_{y}}+B_{z}\overrightarrow{e_{z}};\\ \\ \overrightarrow{v}=v_{x}\overrightarrow{e_{x}}+v_{y}\overrightarrow{e_{y}}+v_{z}\overrightarrow{e_{z}}.\end{split}$ (S5)
Thus, the S4 can be rewritten as:
$\begin{split}\overrightarrow{\tau}=(P_{x}v_{x}+P_{y}v_{y}+P_{z}v_{z})\cdot(B_{x}\overrightarrow{e_{x}}+B_{y}\overrightarrow{e_{y}}+B_{z}\overrightarrow{e_{z}})\\ \\ -(P_{x}B_{x}+P_{y}B_{y}+P_{z}B_{z})\cdot(v_{x}\overrightarrow{e_{x}}+v_{y}\overrightarrow{e_{y}}+v_{z}\overrightarrow{e_{z}})\\ \\ =(P_{y}v_{x}B_{x}+P_{z}v_{z}B_{x}-P_{y}B_{y}v_{x}-P_{z}B_{z}v_{x})\cdot\overrightarrow{e_{x}}\\ \\ +(P_{x}v_{y}B_{y}+P_{z}v_{z}B_{y}-P_{x}B_{x}v_{y}-P_{z}B_{z}v_{y})\cdot\overrightarrow{e_{y}}\\ \\ +(P_{x}v_{x}B_{z}+P_{y}v_{y}B_{z}-P_{x}B_{x}v_{z}-P_{y}B_{y}v_{z})\cdot\overrightarrow{e_{z}}\end{split}$ (S6)
S4 The angular dependence of the torque and the observed bi-axial symmetry
In our experiment, the magnetic field was rotated in the $xz$ plane, the longitudinal thermal gradient thermal gradient was applied along the $x$-axis. Putting $B_{y}=0$ and $P_{y}=0$, Equation S6 becomes:
$\begin{split}\overrightarrow{\tau}=(P_{z}v_{z}B_{x}-P_{z}B_{z}v_{x})\cdot\overrightarrow{e_{x}}\\ \\ -(P_{x}B_{x}v_{y}+P_{z}B_{z}v_{y})\cdot\overrightarrow{e_{y}}\\ \\ +(P_{x}v_{x}B_{z}-P_{x}B_{x}v_{z})\cdot\overrightarrow{e_{z}}\end{split}$ (S7)
We measured the transverse thermal gradient along the $z$-axis. The torque along the z-axis is equal to :
$\tau_{z}=(P_{x}v_{x})B_{z}-(P_{x}v_{z})B_{x}$ (S8)
One can see that the expected torque along the $z$-axis has two components. Each of them vanishes when the field is along either the $x$-axis or the $z$-axis. This is in agreement with the experimental data seen in Fig. 3 of the main text.

### v2

@@
 The long phonon mean-free-path leaves little room for scattering by point-like symmetry breaking defects.
-we show that the signal peaks when the magnetic field is oriented along one of the two diagonal orientations of the puckered honeycomb plane and argue that this can be understood as the sum of two distinct contributions each parallel to a mirror plane.
-This angular dependence as well as the order of magnitude of the observed signal points to the torque exerted by magnetic field on electric dipoles traveling with heat-carrying phonons as the driver of the effect.
+We argue that the angular dependence of the signal can be accounted for as a sum of two sinusoidal components each peaking when the magnetic field is parallel to a high symmetry axis.
+Anharmonicity and finite thermal expansion emerge as indispensable ingredients for a planar Hall signal.
+We identify the torque exerted by magnetic field on traveling electric dipolar waves as a possible microscopic driver.
 The thermal conductivity is a second rank tensor linking the heat current density and the temperature gradient vectors.
 The thermal Hall effect (THE) refers to non-zero off-diagonal components of this tensor, $\overline{\kappa}$.
-Its phonon-driven origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26] interest.
-More recently, attention has focused on the observation of a planar thermal Hall effect [16, 17, 15].
-This is a configuration with the three relevant vectors (Heat current, temperature gradient and magnetic field) lying in the same plane.
-The persistence of the signal when the signal is forbidden by crystal symmetry was attributed to defects breaking the local symmetry [17].
+Its origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26] interest.
+The planar thermal Hall effect refers to a configuration in which the three relevant vectors (heat current, temperature gradient and magnetic field) lie in the same plane.
+This was first observed in the Kitaev spin liquid candidate $\alpha$-RuCl3 [4, 27, 28, 29].
+More recently, it was also observed in other solids, where the thermal Hall signal is attributed to phonons [30, 31, 32].
+The persistence of the signal, when it is forbidden by crystal symmetry, was attributed to defects breaking the local symmetry [31].
 Figure 1: Crystal structure, phonon mean free path and two versions of the thermal Hall effect.(a) Top (two layers) and side (single layer) view of the lattice of black phosphorus (BP).
 P atoms, marked in blue and red, belong to atomic sites with distinct environments.
-The $x$ and $z$ axes respectively correspond to the armchair and zigzag directions of the BP layer plane.
 (b) Longitudinal thermal conductivity (from previous work [15]) along the armchair ($\kappa_{xx}$) and the zigzag ($\kappa_{zz}$) orientation, at least four times of anisotropy can be seen.
-(c) The temperature dependence of mean free path of phonon ($l_{ph}$) in BP for two in-plane orientations ($J_{q}||z$ axis and $J_{q}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat.
+(c) The temperature dependence of mean free path of phonon ($\ell_{ph}$) in BP for two in-plane orientations ($J_{q}||z$ axis and $J_{q}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat.
 It shows an increasing behavior before 10 K with cooling and is comparable to the sample thickness around 20 K, the peak temperature of $\kappa_{ii}$.
@@
 Figure 2: Planar thermal Hall data.
-(a) The field dependence of thermal Hall angle $\nabla T_{j}/\nabla T_{i}$ , with $J_{q}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures.
+(a) The field dependence of thermal Hall angle $\nabla_{j}T/\nabla_{i}T$ , with $J_{q}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures.
 This allows to quantify $\kappa_{xz}^{x}$.
@@
 All measurements are performed on the same sample.
-Note that the thermal Hall angle, $\nabla T_{j}/\nabla T_{i}$ is almost unchanged when the orientations of $B$ rotates from $J_{q}$ to $(\nabla T)_{\perp}$.
+Note that the thermal Hall angle, $\nabla_{j}T/\nabla_{i}T$ is almost unchanged when the orientations of $B$ rotates from $J_{q}$ to $(\nabla T)_{\perp}$.
 On the other hand, it changes by a factor of six times when the orientations of $J_{q}$ and $(\nabla T)_{\perp}$ permute with each other.
@@
 […]
-Figure 4: From atomic displacements to dipolar moments (a) Lattice distortion due to the thermal expansion.
+Figure 4: From atomic displacements to dipolar moments (a) Because of thermal expansion, the orthorhombic unit cell is distorted.
+A rectangle with two perpendicular mirrors is replaced by a trapezoid with no reflection symmetry.
 (b) Phosphorus atoms surrounding an inversion center.
@@
 (d) This torque ($\tau=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$) will survive only when the magnetic field is oriented perpendicular to the wave velocity.
-Orange arrows show the orientation of the Lorenz forces on the two poles of the dipole.
-Here, we report on the observation of a planar Hall effect in black phosphorus(P), an elemental insulator with ballistic phonons [27] and no magnetism.
+Green arrows show the orientation of the Lorenz forces on the two poles of the dipole.
+Here, we report on the observation of a planar Hall effect in black phosphorus(P), an elemental insulator with ballistic phonons [33].
 The thermal conductivity tensor has unequal diagonal components ($\kappa_{zz}\simeq 5\kappa_{xx}$).
@@
 A quantitative account of our observation is missing.
-Nevertheless, we argue that a scenario in which electric dipole moments, generated by the thermal gradient and the atomic vibrations, play a key role is compatible with the order of magnitude of the observed signal as well as its angular variation.
+Nevertheless, we argue that because of the finite thermal expansion, a temperature gradient allows the emergence of a signal forbidden in thermodynamic equilibrium.
+Electric dipole moments travelling with sound can couple to magnetic field and generate a thermal Hall signal in planar configuration.
 Fig.
@@
 2a-c shows the planar thermal Hall data for three different configurations.
-In each panel, the orientations of the three relevant vectors (B, $J_{q}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla T_{j}/\nabla T_{i}$) is shown.
+In each panel, the orientations of the three relevant vectors ($B$, $J_{q}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla_{j}T/\nabla_{i}T$) is shown.
 Fig.
@@
 1c) are both anisotropic at zero magnetic field.
-The issue is to find a way for magnetic field to skew one or both of these tensors off the symmetry axes
+The issue is to find a way for magnetic field to skew one or both of these tensors off the symmetry axes.
 Recent theoretical studies have scrutinized intrinsic phonon band[37] incorporated with energy magnetization contribution [38, 21] and phonon angular momentum acquired through interaction with the spins [39, 40] or with orbital motion of ions [41].
@@
 Nevertheless, any phonon breaking the inversion symmetry will generate an electric dipolar wave.
-This is a feature highlight in the context of ferrons, the elementary excitation of a ferroelectric solid, even in a paraelectric insulators [38, 39].
-The thermal expansion of black P is not only anisotropic, but has opposite signs along armchair and zigzag crystalline orientations [40].
-As a consequence, in presence of a temperature gradient each tetragonal unit cell is slightly distorted.
-Its lattice parameter, compared to its colder or warmer neighboring cells, is longer and shorter along opposite (Fig.
-4a) orientations.
-This paves the way for finite dipolar polarization, and for chirality in presence of the magnetic field.
+This feature highlighted in the context of ferrons, the elementary excitation of a ferroelectric solid [44, 45], persists even in a paraelectric insulator.
+Black P has an anisotropic thermal expansion with opposite signs along armchair and zigzag crystalline orientations [46].
+In presence of a temperature gradient, each unit cell is slightly distorted (Fig.
+4a).
+This feature, driven by unavoidable anharmonicity in any solid [47], is the fundamental reason that a signal forbidden in thermodynamic equilibrium is allowed in presence of a temperature gradient.
+It also paves the way for chirality in presence of the magnetic field.
 The inversion center of the pristine unit cell is not an atomic site and atomic displacements associated with phonons generate electric dipole moments (Fig.
@@
 However, let us compare the order of magnitude of the expected and the measured signals.
-Experimentally, the thermal Hall angle at $B$ = 10 T peaks to ×$\approx\times 10^{-2}$ in black P.
+Experimentally, the thermal Hall angle at $B$ = 10 T peaks to $\approx 10^{-2}$ in black P.
 The length scale extracted from this angle and fundamental constants ($\lambda_{tha}$ = $\ell_{B}\sqrt{\kappa_{ij}/\kappa_{jj}}$) is about 5 Å.
 Intriguingly, in all insulators displaying a thermal Hall signal, this length scale lies in a narrow range (2 Å${}<\lambda_{tha}<7$ [15, 17].
-As seen in figure S2 in the supplement [30], this remains true for recently studied cases of Si, Ge and quartz [15].
+As seen in Fig.
+S2 in the supplement [36], this remains true for recently studied cases of Si, Ge and quartz [32].
 Suppose an electric dipole moment of $\delta p$ traveling with the velocity of sound, $v_{s}$ (loosely linked to the Debye frequency by $v_{s}\sim a\omega_{D}$, where $a$ is the interatomic distance).
@@
 $\theta_{H}\approx\frac{a^{2}}{\ell_{B}^{2}}$ (5)
-This expectation is of the order of magnitude of the measured signal.
-A rigorous treatment would presumably include in this expression the Grüneisen parameter [42].
-Possibly, the atomic electric dipole polarizability [43] would also be present.
-However, the former, which is dimensionless remains of the order of unity and the latter introduces a length scale which does not vary widely among different solids.
-Therefore, a scenario attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves emerges promising from our study.
+This is of the order of magnitude of the measured signal.
+A rigorous treatment would presumably include the Grüneisen parameter [47] and the atomic electric dipole polarizability.
+The former, which is dimensionless, remains of the order of unity and the latter introduces a length scale which does not vary widely among different solids [49].
+Thus, a scenario attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves, which to the best of our knowledge, was not considered before, emerges as a promising candidate.
+Acknowledgements.
 We thank Benoît Fauqué for stimulating discussions.
@@
 12304065).
-Supplementary Materials for “Angle-dependent planar thermal Hall effect in black phosphorus”
+Supplementary Materials for “Angle-dependent planar thermal Hall effect by quasi-ballistic phonons in black phosphorus” by X. Li et al.
 S1 Materials and Methods
@@
 […]
-The longitudinal ($\nabla T_{i}$) and the transverse ($\nabla T_{j}$ ) thermal gradient generated by a longitudinal thermal current $J_{q}$ were measured.
-They lead to the longitudinal ($\kappa_{ii}$) and the transverse ($\kappa_{ij}$) thermal conductivity, as well as the thermal Hall angle ($\nabla T_{j}/\nabla T_{i}$):
-$\kappa_{ii}=\frac{Q_{i}}{\nabla T_{i}}$ (S1)
-$\frac{\nabla T_{j}}{\nabla T_{i}}=\frac{\kappa_{ij}}{\kappa_{jj}}$ (S2)
-$\kappa_{ij}=\frac{\nabla T_{j}}{\nabla T_{i}}\cdot\kappa_{jj}$ (S3)
+The longitudinal ($\nabla_{i}T$) and the transverse ($\nabla_{j}T$) thermal gradient generated by a longitudinal thermal current $J_{q}$ were measured.
+They lead to the longitudinal ($\kappa_{ii}$) and the transverse ($\kappa_{ij}$) thermal conductivity, as well as the thermal Hall angle ($\nabla_{j}T/\nabla_{i}T$):
+$\kappa_{ii}=\frac{Q_{i}}{\nabla_{i}T}$ (S1)
+$\frac{\nabla_{j}T}{\nabla_{i}T}=\frac{\kappa_{ij}}{\kappa_{jj}}$ (S2)
+$\kappa_{ij}=\frac{\nabla_{j}T}{\nabla_{i}T}\cdot\kappa_{jj}$ (S3)
 Here $Q$ is the heat power.
@@
 Fig.
-S1 shows the field dependent thermal Hall angles ($\nabla T_{j}/\nabla T_{i}$) with the $J_{q}$ alongs $x$ axis and the $(\nabla T)_{\perp}$ along $z$ axis at six different field orientations.
+S1 shows the field dependent thermal Hall angles ($\nabla_{j}T/\nabla_{i}T$) with the $J_{q}$ alongs $x$ axis and the $(\nabla T)_{\perp}$ along $z$ axis at six different field orientations.
 $\phi$ is the angle of the magnetic field with respect to the $J_{q}$ ($x$ axis).
-$\phi=0^{\circ}$ and $\phi=90^{\circ}$ represent the orientations along two high-symmetry axes, and their $\nabla T_{j}/\nabla T_{i}$ signal have the same amplitude but opposite signs.
+$\phi=0^{\circ}$ and $\phi=90^{\circ}$ represent the orientations along two high-symmetry axes, and their $\nabla_{j}T/\nabla_{i}T$ signal have the same amplitude but opposite signs.
 $\phi=-45^{\circ}$ and $\phi=45^{\circ}$ represent two diagonal orientations of the puckered honeycomb plane that have the maximum and vanished responses respectively, suggesting that the planar thermal Hall signal consists of two distinct contributions each along one high-symmetry axis.
@@
 Figure S1: Raw data of angular dependent planar thermal Hall effect in black phosphorus.
-Field dependent thermal Hall angle ($\nabla T_{z}/\nabla T_{x}$) at a constant temperature and six different field orientations.
-Figure S2: Thermal Hall angle in different insulators The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source:  [1, 2, 3, 4, 5, 6, 8, 9, 11, 12, 14, 15]).
+Field dependent thermal Hall angle ($\nabla_{z}T/\nabla_{x}T$) at a constant temperature and six different field orientations.
+Figure S2: Thermal Hall angle in different insulators.
+The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source: [1, 50, 6, 7, 8, 9, 11, 12, 14, 15, 17, 32]).
+Note that magnetic insulators with a non-linear field dependence of the thermal Hall effect, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.
 S3 Torque exerted by the magnetic field on moving dipoles
@@
 S4 The angular dependence of the torque and the observed bi-axial symmetry
-In our experiment, the magnetic field was rotated in the $xz$ plane, the longitudinal thermal gradient thermal gradient was applied along the $x$-axis.
-Putting $B_{y}=0$ and $P_{y}=0$, Equation S6 becomes:
+In our experiment, the magnetic field was rotated in the $xz$ plane, the longitudinal thermal gradient was applied along the $x$-axis. Putting $B_{y}=0$ and $P_{y}=0$, Equation S6 becomes:
 $\begin{split}\overrightarrow{\tau}=(P_{z}v_{z}B_{x}-P_{z}B_{z}v_{x})\cdot\overrightarrow{e_{x}}\\ \\ -(P_{x}B_{x}v_{y}+P_{z}B_{z}v_{y})\cdot\overrightarrow{e_{y}}\\ \\ +(P_{x}v_{x}B_{z}-P_{x}B_{x}v_{z})\cdot\overrightarrow{e_{z}}\end{split}$ (S7)
@@
 Each of them vanishes when the field is along either the $x$-axis or the $z$-axis.
-This is in agreement with the experimental data seen in Fig.
-3 of the main text.
+This is in agreement with the experimental data seen in Figure 3 of the main text.
+S5 The universality in thermal Hall physics
+Figure S2 illustrates the thermal Hall angle across various insulators.
+It presents the ratio of transverse thermal conductivity to magnetic field as a function of longitudinal thermal conductivity for these insulators.
+It is important to note that the figure exclusively features insulators exhibiting a linear or sub-linear relationship between transverse thermal conductivity and magnetic field.
+Magnetic insulators with a non-linear dependence, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.

### v3

@@
 The origin of the phonon thermal Hall effect in insulators is a matter of ongoing debate.
-The large amplitude of the signal in an elemental non-magnetic solid, such as black phosphorus (BP) calls for a minimal mechanism with no role for spin degree of freedom.
+The large amplitude of the signal in an elemental non-magnetic solid, such as black phosphorus (BP), calls for a minimal mechanism not invoking the spin degree of freedom.
 Here, we show that a longitudinal heat flow generates a transverse temperature gradient in BP even when the magnetic field, the heat current and the thermal gradient lie in the same plane.
-The long phonon mean-free-path leaves little room for scattering by point-like symmetry breaking defects.
-We argue that the angular dependence of the signal can be accounted for as a sum of two sinusoidal components each peaking when the magnetic field is parallel to a high symmetry axis.
-Anharmonicity and finite thermal expansion emerge as indispensable ingredients for a planar Hall signal.
-We identify the torque exerted by magnetic field on traveling electric dipolar waves as a possible microscopic driver.
+The phonon mean-free-path is close to the sample thickness.
+Therefore, it is unlikely that scattering by point-like symmetry-breaking defects play a major role.
+We show that the angular dependence of the signal can be mapped to the sum of two sinusoidal components each peaking when the magnetic field is parallel to a high symmetry.
+We propose that anharmonicity may play a major role and argue that the magnetic field can exert a torque on electric dipolar waves traveling with phonons.
+Keywords: Phonon, Planar thermal Hall effect, Black phosphorus, Anharmonicity
+I 1. Introduction
 The thermal conductivity is a second rank tensor linking the heat current density and the temperature gradient vectors.
 The thermal Hall effect (THE) refers to non-zero off-diagonal components of this tensor, $\overline{\kappa}$.
-Its origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26] interest.
+Its origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26, 27] interest.
 The planar thermal Hall effect refers to a configuration in which the three relevant vectors (heat current, temperature gradient and magnetic field) lie in the same plane.
@@
 (b) Longitudinal thermal conductivity (from previous work [15]) along the armchair ($\kappa_{xx}$) and the zigzag ($\kappa_{zz}$) orientation, at least four times of anisotropy can be seen.
-(c) The temperature dependence of mean free path of phonon ($\ell_{ph}$) in BP for two in-plane orientations ($J_{q}||z$ axis and $J_{q}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat.
+(c) The temperature dependence of mean free path of phonon ($\ell_{\mathrm{ph}}$) in BP for two in-plane orientations ($J_{\mathrm{q}}||z$ axis and $J_{\mathrm{q}}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat.
 It shows an increasing behavior before 10 K with cooling and is comparable to the sample thickness around 20 K, the peak temperature of $\kappa_{ii}$.
-(d) The schematic for measuring the thermal Hall effect, with the magnetic field ($B$) perpendicular to the longitudinal heat current ($J_{q}$) and the transverse thermal gradient ($(\nabla T)_{\perp}$).
+(d) The schematic for measuring the thermal Hall effect, with the magnetic field ($B$) perpendicular to the longitudinal heat current ($J_{\mathrm{q}}$) and the transverse thermal gradient ($(\nabla T)_{\perp}$).
 The thermal Hall conductivity tensor is denoted as $\kappa_{ij}^{k}$.
-(e) The schematic for measuring planar thermal Hall effect, with the magnetic field ($B$) rotating from the orthogonal to the parallel direction of the $J_{q}$ or the $(\nabla T)_{\perp}$.
+(e) The schematic for measuring planar thermal Hall effect, with the magnetic field ($B$) rotating from the orthogonal to the parallel direction of the $J_{\mathrm{q}}$ or the $(\nabla T)_{\perp}$.
 Two types of planar thermal Hall conductivity tensors are denoted as $\kappa_{ij}^{i}$ and $\kappa_{ij}^{j}$ respectively.
 Figure 2: Planar thermal Hall data.
-(a) The field dependence of thermal Hall angle $\nabla_{j}T/\nabla_{i}T$ , with $J_{q}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures.
+(a) The field dependence of thermal Hall angle $\nabla_{j}T/\nabla_{i}T$ , with $J_{\mathrm{q}}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures.
 This allows to quantify $\kappa_{xz}^{x}$.
-(b) Same for $J_{q}$ along the $x$ axis, $B$ and $(\nabla T)_{\perp}$ along the $z$ axis.
+(b) Same for $J_{\mathrm{q}}$ along the $x$ axis, $B$ and $(\nabla T)_{\perp}$ along the $z$ axis.
 This allows to quantify $\kappa_{xz}^{z}$.
-(c) Same for $J_{q}$ and $B$ along the $z$ axis and $(\nabla T)_{\perp}$ along the $x$ axis.
+(c) Same for $J_{\mathrm{q}}$ and $B$ along the $z$ axis and $(\nabla T)_{\perp}$ along the $x$ axis.
 This allows to quantify $\kappa_{zx}^{z}$.
@@
 All measurements are performed on the same sample.
-Note that the thermal Hall angle, $\nabla_{j}T/\nabla_{i}T$ is almost unchanged when the orientations of $B$ rotates from $J_{q}$ to $(\nabla T)_{\perp}$.
-On the other hand, it changes by a factor of six times when the orientations of $J_{q}$ and $(\nabla T)_{\perp}$ permute with each other.
+Note that the thermal Hall angle, $\nabla_{j}T/\nabla_{i}T$ is almost unchanged when the orientations of $B$ rotates from $J_{\mathrm{q}}$ to $(\nabla T)_{\perp}$.
+On the other hand, it changes by a factor of six times when the orientations of $J_{\mathrm{q}}$ and $(\nabla T)_{\perp}$ permute with each other.
 This difference combined with the anisotropy in longitudinal thermal conductivity warrants Onsager reciprocity: $\kappa_{xz}^{z}=-\kappa_{zx}^{z}$, as seen in (e).
-(f) Comparison of thermal Hall conductivity $\kappa_{xz}$ with the magnetic field $B$ rotating from the orthogonal ($y$ axis, from previous work [15] with the same sample) to the parallel direction of the $J_{q}$ ($x$ axis) or the $(\nabla T)_{\perp}$ ($z$ axis).
+(f) Comparison of thermal Hall conductivity $\kappa_{xz}$ with the magnetic field $B$ rotating from the orthogonal ($y$ axis, from previous work [15] with the same sample) to the parallel direction of the $J_{\mathrm{q}}$ ($x$ axis) or the $(\nabla T)_{\perp}$ ($z$ axis).
 […]
-Figure 4: From atomic displacements to dipolar moments (a) Because of thermal expansion, the orthorhombic unit cell is distorted.
-A rectangle with two perpendicular mirrors is replaced by a trapezoid with no reflection symmetry.
-(b) Phosphorus atoms surrounding an inversion center.
+Figure 4: Unit cell in presence of a heat current and magnetic field.
+(a) The pristine unit cell without the heat current and the magnetic field.
+(b) The unit cell in presence of a heat current $J_{\mathrm{q}}$, It changed from a rectangle to a trapezoid.
+The only mirror plane which survives is the one parallel to the heat current.
+Which vanishes too in presence of a magnetic field $B$, as seen in (c).
+The mirror planes not withstanding, a planar Hall response can emerge.
+(d) Changing the orientation of the heat current will change the way the primitive cell is distorted.
+In this case, applying a magnetic field along the same orientation will generate a transverse thermal gradient of opposite sign (e).
+On the other hand, inverting the orientation of the magnetic field will double flip the orientation of the transverse thermal gradient (f).
+Figure 5: From atomic displacements to dipolar moments.
+(a) Phosphorus atoms surrounding an inversion center.
 Their displacements associated with an acoustic mode (top) will affect the local electric dipole moments (middle), generating a net dipole moment (bottom).
-(c) Phonon modes are associated with waves of dipolar moments.
+(b) Phonon modes are associated with waves of dipolar moments.
 A static magnetic field will exert a torque on the dipole moment traveling with the sound wave.
-(d) This torque ($\tau=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$) will survive only when the magnetic field is oriented perpendicular to the wave velocity.
+(c) This torque ($\tau=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$) will survive only when the magnetic field is oriented perpendicular to the wave velocity.
 Green arrows show the orientation of the Lorenz forces on the two poles of the dipole.
@@
 Electric dipole moments travelling with sound can couple to magnetic field and generate a thermal Hall signal in planar configuration.
+II 2. Materials and methods
+Detailed materials and methods can be found in the supplementary materials, including sample details, thermal transport measurement methods.
+III 3. Results
 Fig.
@@
 4.6 km/s along the $x$-axis) [36].
-As a consequence, the phonon mean free path $\ell_{ph}(T)$ , shown in Fig.
+As a consequence, the phonon mean free path $\ell_{\mathrm{ph}}(T)$, shown in Fig.
 1c, is also anisotropic.
@@
 A Hall response refers to a signal odd in magnetic field and oriented perpendicular to the applied current.
-Designating the transverse temperature gradient by $(\nabla T)_{\perp}$, the applied longitudinal heat current by $J_{q}$ and the magnetic field by $B$, the standard thermal Hall configuration corresponds to $B\perp J_{q}\perp(\nabla T)_{\perp}$.
+Designating the transverse temperature gradient by $(\nabla T)_{\perp}$, the applied longitudinal heat current by $J_{\mathrm{q}}$ and the magnetic field by $B$, the standard thermal Hall configuration corresponds to $B\perp J_{\mathrm{q}}\perp(\nabla T)_{\perp}$.
 It is sketched in panel d and was the subject of our previous study [15].
-Panel e shows the configuration for planar thermal Hall effect, which corresponds to $B~{}||~{}(J_{q}\perp(\nabla T)_{\perp})$.
-We use $\kappa_{ij}^{k}$ to designate the $\overline{\kappa}$ component corresponding to $J_{q}$ along the $i$-axis, $(\nabla T)_{\perp}$ along the $j$-axis and the magnetic field along the $k$-axis.
+Panel e shows the configuration for planar thermal Hall effect, which corresponds to $B~{}||~{}(J_{\mathrm{q}}\perp(\nabla T)_{\perp})$.
+We use $\kappa_{ij}^{k}$ to designate the $\overline{\kappa}$ component corresponding to $J_{\mathrm{q}}$ along the $i$-axis, $(\nabla T)_{\perp}$ along the $j$-axis and the magnetic field along the $k$-axis.
 Fig.
-2a-c shows the planar thermal Hall data for three different configurations.
-In each panel, the orientations of the three relevant vectors ($B$, $J_{q}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla_{j}T/\nabla_{i}T$) is shown.
+2a–c shows the planar thermal Hall data for three different configurations.
+In each panel, the orientations of the three relevant vectors ($B$, $J_{\mathrm{q}}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla_{j}T/\nabla_{i}T$) is shown.
 Fig.
 2d compares the temperature dependence of the thermal Hall angle in three planar configurations at 12 T.
-It is finite in the three planar configurations, but there is a large difference in amplitude when one permutes the orientations of $J_{q}$ and $(\nabla T)_{\perp}$.
+It is finite in the three planar configurations, but there is a large difference in amplitude when one permutes the orientations of $J_{\mathrm{q}}$ and $(\nabla T)_{\perp}$.
 Since the longitudinal thermal conductivity is anisotropic by the same factor, this difference warrants an equality between the absolute values of $\kappa_{xz}^{z}$ and $\kappa_{zx}^{z}$, as expected by Onsager reciprocity.
@@
 2f).
+We then proceeded to quantify the angle dependence of this planar Hall signal at fixed temperature as the magnetic field rotated in the $xz$ plane.
+The raw data, shown in the Supplementary material, shows a field-linear response for all explored angles.
+The measurements were performed at 44.5 K.This temperature was chosen to strike a balance between proximity to the peak temperature and the optimal sensitivity of the thermocouples.
+As seen in Fig.
+3a, the signal exhibits a twofold oscillation.
+It peaks (with opposite signs) at $\phi=-\pi/4$ and $\phi=3\pi/4$, i.e., along one of the diagonals of the $xz$ plane, and vanishes along the other diagonal.
 […]
+IV 4. Discussion and conclusion
 Thus, we detect in a crystal belonging with the $mmm$ point group symmetry, there is a planar thermal Hall signal with bi-axial symmetry.
-Let us note that at peak temperature of this thermal Hall signal, the longitudinal thermal conductivity of black P is as large as $\kappa_{p}\simeq 2000$ W/Km.
-This is to be compared with cases such as $\alpha$-RuCl3 ($\kappa_{p}\simeq$ 3 W/Km) [14], Na2Co2TeO6 ($\kappa_{p}\simeq$ 10 W/Km) [30] or Y-kapellasite ($\kappa_{p}\simeq$ 2 W/Km) [31].
-Contrary to these cases, the phonon mean free path in black P is close to the sample dimensions and local defects are unlikely to play a major role.
+Let us note that at peak temperature of this thermal Hall signal, the longitudinal thermal conductivity of black P is as large as $\kappa_{\mathrm{p}}\simeq 2000$ WK-1m-1.
+This is to be compared with cases such as $\alpha$-RuCl3 ($\kappa_{\mathrm{p}}\simeq$ 3 WK-1m-1) [14], Na2Co2TeO6 ($\kappa_{\mathrm{p}}\simeq$ 10 WK-1m-1) [31] or Y-kapellasite ($\kappa_{\mathrm{p}}\simeq$ 2 WK-1m-1) [32].
+The thermal conductivity of black phosphorus is three orders of magnitude larger, with a phonon mean free path close to the sample thickness.
+These features strongly suggest that the thermal transport mechanism in black phosphorus is dominated by boundary scattering and phonon-phonon scattering, rather than being influenced by local defect scattering.
+However, we cannot totally exclude a minor role played by local defects.
 In anisotropic dense medium, thermal diffusivity is a tensor, $\overline{D}$. The heat equation becomes:
@@
 Consider now the energy continuity equation, which relates the energy flux to specific heat per volume, $C$:
-$\overrightarrow{\nabla}\cdot\overrightarrow{J_{q}}+C\frac{\partial T}{\partial t}=0$ (2)
+$\overrightarrow{\nabla}\cdot\overrightarrow{J_{\mathrm{q}}}+C\frac{\partial T}{\partial t}=0$ (2)
 The combination of the two previous equations leads to:
-$\overrightarrow{J_{q}}=-C\overline{D}\overrightarrow{\nabla}T$ (3)
+$\overrightarrow{J_{\mathrm{q}}}=-C\overline{D}\overrightarrow{\nabla}T$ (3)
 The thermal conductivity tensor $\overline{\kappa}$, is thus the thermal diffusivity tensor multiplied by a scalar: $\overline{\kappa}=\overline{D}C$.
@@
 The latter is the product of carrier velocity and carrier mean free path.
-Sound velocity, $v_{s}$, and phonon mean free path $\ell_{ph}$ (see Fig.
+Sound velocity, $v_{\mathrm{s}}$, and phonon mean free path $\ell_{\mathrm{ph}}$ (see Fig.
 1c) are both anisotropic at zero magnetic field.
@@
 These are optical phonons and not heat-carrying acoustic modes which interest us.
+The lowest optical phonon frequency of black phosphorus is around 140 cm-1 [43] (see the Supplementary material for more details), much larger than the acoustic phonons.
+However, for the high temperature signals, the indirect contribution of optical phonons through coupling with acoustic phonons may not be negligible.
 Nevertheless, any phonon breaking the inversion symmetry will generate an electric dipolar wave.
@@
 Black P has an anisotropic thermal expansion with opposite signs along armchair and zigzag crystalline orientations [46].
-In presence of a temperature gradient, each unit cell is slightly distorted (Fig.
-4a).
+In presence of a temperature gradient, each unit cell is slightly distorted (See Fig.
+4a, b and d).
 This feature, driven by unavoidable anharmonicity in any solid [47], is the fundamental reason that a signal forbidden in thermodynamic equilibrium is allowed in presence of a temperature gradient.
-It also paves the way for chirality in presence of the magnetic field.
+It also paves the way for chirality in presence of the magnetic field (See Fig.
+4c, e and f).
+Let us now draw a tentative picture of how the magnetic field may couple to heat-carrying phonons.
 The inversion center of the pristine unit cell is not an atomic site and atomic displacements associated with phonons generate electric dipole moments (Fig.
-4b).
+5a).
 Therefore, acoustic phonons breaking inversion centers are traveling electric dipolar moments capable of coupling to a static magnetic field (Fig.
-4c).
+5b).
+Electric dipole wave and its evolution under magnetic field can be detected via methods such as $P$-$E$ response and second harmonic generation directly in the future.
 The torque exerted by a static magnetic field, $\overrightarrow{B}$, on an electric dipole of $\overrightarrow{P}$ moving with a velocity of $\overrightarrow{v}$ is : $\overrightarrow{\tau}=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$ [48].
@@
 Thus, the torque due to the Lorenz force exerted on each of the two poles is finite when the magnetic field is oriented perpendicular to the orientation of the dipole propagation (Fig.
-4d).
-These are ingredients for a scenario for planar thermal Hall signal in a non-magnetic insulator (see the supplement [36] for a discussion of the expected angular variation).
-A rigorous treatment is beyond the scope of this paper.
+5c).
+These are ingredients for a scenario for planar thermal Hall signal in a non-magnetic insulator (see the Supplementary material for a discussion of the expected angular variation).
+A rigorous treatment is beyond the scope of this paper, but torque is known to be responsible for changes in angular momentum, so a quantitative theoretical analysis could involve either the phonon angular momentum [39, 49] or the phonon Berry curvature [20].
 However, let us compare the order of magnitude of the expected and the measured signals.
 Experimentally, the thermal Hall angle at $B$ = 10 T peaks to $\approx 10^{-2}$ in black P.
-The length scale extracted from this angle and fundamental constants ($\lambda_{tha}$ = $\ell_{B}\sqrt{\kappa_{ij}/\kappa_{jj}}$) is about 5 Å.
-Intriguingly, in all insulators displaying a thermal Hall signal, this length scale lies in a narrow range (2 Å${}<\lambda_{tha}<7$ [15, 17].
+The length scale extracted from this angle and fundamental constants ($\lambda_{\mathrm{tha}}$ = $\ell_{B}\sqrt{\kappa_{ij}/\kappa_{jj}}$) is about 5 Å.
+Intriguingly, in all insulators displaying a thermal Hall signal, this length scale lies in a narrow range  [15, 17] (2 Å${}<\lambda_{\mathrm{tha}}<7$).
 As seen in Fig.
-S2 in the supplement [36], this remains true for recently studied cases of Si, Ge and quartz [32].
-Suppose an electric dipole moment of $\delta p$ traveling with the velocity of sound, $v_{s}$ (loosely linked to the Debye frequency by $v_{s}\sim a\omega_{D}$, where $a$ is the interatomic distance).
+6, this remains true for recently studied cases of Si, Ge and quartz [33].
+Figure 6: Thermal Hall angle in different insulators.
+The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source: [1, 50, 6, 7, 8, 9, 11, 12, 14, 15, 17, 33]).
+Note that magnetic insulators with a non-linear field dependence of the thermal Hall effect, such as Lu2V2O7[2] and Cu(1-3, bdc)[51], are not included in the figure.
+Suppose an electric dipole moment of $\delta p$ traveling with the velocity of sound, $v_{\mathrm{s}}$ (loosely linked to the Debye frequency by $v_{\mathrm{s}}\sim a\omega_{\mathrm{D}}$, where $a$ is the interatomic distance).
 The exerted torque, that is the angular derivative of the energy, would be $Ba\omega_{D}\delta p$.
@@
 The former, which is dimensionless, remains of the order of unity and the latter introduces a length scale which does not vary widely among different solids [52].
-Thus, a scenario attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves, which to the best of our knowledge, was not considered before, emerges as a promising candidate.
-Acknowledgements.
+This admittedly hand-waving scenario, attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves, was not considered before.
+Moreover, its influence merits in-depth consideration when applied to other elemental insulators.
+V Conflict of interest
+The authors declare that they have no conflict of interest.
+VI Acknowledgments
 We thank Benoît Fauqué for stimulating discussions.
-This work was supported by the National Key Research and Development Program of China (Grant No.
-2022YFA1403500), the National Science Foundation of China (Grant No.
-12004123, 51861135104 and 11574097 ), and the Fundamental Research Funds for the Central Universities (Grant no.
-2019kfyXMBZ071).
+This work was supported by the National Key Research and Development Program of China (2022YFA1403500), the National Natural Science Foundation of China (12004123, 51861135104 and 11574097 ), and the Fundamental Research Funds for the Central Universities (2019kfyXMBZ071).
 X.
 L.
-was supported by The National Key Research and Development Program of China (Grant No.2023YFA1609600) and the National Science Foundation of China (Grant No.
-12304065).
+was supported by The National Key Research and Development Program of China (2023YFA1609600) and the National Natural Science Foundation of China (12304065).
+VII Author contributions
+Xiaokang Li, Zengwei Zhu, and Kamran Behnia proposed and supervised the project.
+Xiaokang Li and Xiaodong Guo prepared the samples, designed and performed the measurements.
+Xiaokang Li, Zengwei Zhu, and Kamran Behnia analyzed data and wrote the manuscript with the contributions of all authors.
 Supplementary Materials for “Angle-dependent planar thermal Hall effect by quasi-ballistic phonons in black phosphorus” by X. Li et al.
@@
 Field dependent thermal Hall angle ($\nabla_{z}T/\nabla_{x}T$) at a constant temperature and six different field orientations.
-Figure S2: Thermal Hall angle in different insulators.
-The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source: [1, 50, 6, 7, 8, 9, 11, 12, 14, 15, 17, 32]).
-Note that magnetic insulators with a non-linear field dependence of the thermal Hall effect, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.
 S3 Torque exerted by the magnetic field on moving dipoles
@@
 This is in agreement with the experimental data seen in Figure 3 of the main text.
-S5 The universality in thermal Hall physics
-Figure S2 illustrates the thermal Hall angle across various insulators.
-It presents the ratio of transverse thermal conductivity to magnetic field as a function of longitudinal thermal conductivity for these insulators.
-It is important to note that the figure exclusively features insulators exhibiting a linear or sub-linear relationship between transverse thermal conductivity and magnetic field.
-Magnetic insulators with a non-linear dependence, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.
+Figure S2: Phonon spectrum in black phosphorus. Adapted from [34].
+Mode Calc. Machida et al. Exp. Fujii et al.
+$v_{x}$ LAx 0.454 0.46
+TAz 0.490 0.46
+TAy 0.119
+Average 0.354
+$v_{z}$ LAz 0.833 0.96
+TAx 0.488
+TAy 0.287
+Average 0.536
+$v_{y}$ LAy 0.505 0.510
+TAx 0.160 0.170
+TAz 0.289 0.310
+Average 0.318 0.330
+Table S1: Phonon velocities in black P. Velocities of the three acoustic phonon modes (in units of 104 m/s) along different high symmetry axes
+S5 Phonon velocities and phonon spectrum in black Phosphorus.
+Table S1 lists the velocity of longitudinal (LA) and transverse acoustic (TA) modes in black P according to the calculated phonon spectrum [34] and the neutron scattering experiments [53].
+Note the good agreement between theory and experiment.
+Remarkably, the longitudinal mode is faster than the transverse mode.
+Fig.
+S2 shows the phonon spectrum up to 500 cm-1.
+The lowest optical phonon frequency of black phosphorus is around 140 cm-1, much larger than the acoustic phonons below 40 cm-1 which are the focus of this work.